OpenStudy (anonymous):

Prove the identity(Pre-Calc/Math 4): Sin4x=(4sin(x)cos(x))(2cos^2(x)-1)

OpenStudy (thomas5267):

Hint: \[ \sin(4x)=\sin(2\times 2x) \]

OpenStudy (anonymous):

\[\sin 2a=2\sin a \cos a\] \[\cos 2a=\cos ^2a-\sin ^2a=\cos ^2a-(1-\cos ^2a)=2\cos ^2a-1\]